When working in three-sixteenths inch scale, a standard 8.5" x 11" sheet of paper represents a rectangle with approximate dimensions of how many inches?
Answer by TimothyLamb(4379)

x = real world dimension
y = model dimension
---
y/x = 3/16
x = (16/3)y
---
model width (y) = 8.5:
x = (16/3)y
x = (16/3)*8.5
x = 45.333333333
---
model length (y) = 11:
x = (16/3)y
x = (16/3)*11
x = 58.666666666
---
answer:
w = 45.3 inches
L = 58.7 inches
---
Free algebra tutoring live chat:
https://sooeet.com/chat.php?gn=algebra
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations with quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php
---